# An Efficient Image Denoising Method Integrating Multi-resolution Local Clustering and Adaptive Smoothing

### v2

@@ Abstract
The importance of developing efficient image denoising methods is immense especially for modern applications such as image comparisons, image monitoring, medical image diagnostics, and so forth. Available methods in the vast literature on image denoising can address certain issues in image denoising, but no one single method can solve all such issues. For example, jump regression based methods can preserve linear edges well, but cannot preserve many other fine details of an image. On the other hand, local clustering based methods can preserve fine edge structures, but cannot perform well in presence of heavy noise. The proposed method uses various shapes and sizes of local neighborhood based on local information, and integrates this adaptive approach with the local clustering based smoothing. Theoretical justifications and numerical studies show that the proposed method indeed performs better than these two individual methods and outperforms many other state-of-the-art techniques as well. Such performance demonstrates vast potential of the applicability of the proposed method in many modern-day applications.
Keywords: Adaptive smoothing, Edge preservation, Image denoising, Local clustering, Multi-resolution smoothing, Variable neighborhood.
@@ 1 Introduction
Easy access to modern image acquisition techniques is making images a popular data format in various disciplines of science. In manufacturing industries, applications include stress and strain analysis of products, anomaly detection in the rolling process, steel and tile surface monitoring, etc. In medical science, various imaging modalities such as X-ray, CT-scan, MRI, and fMRI are being widely used for medical diagnosis. For surveillance of the Earth’s surface, satellite images are commonly used. It becomes the basic tool for studying agriculture, forest science, ecology, ecosystems, and many more. Note that in all of these applications, image comparisons and sequential image monitoring are important. These images often contain noise, anomalies, etc. that make eventual image analysis unreliable. Therefore, image analysis as a modern-day research area is gaining a lot of importance. Along with increasing popularity of digital imaging systems, there is a growing demand for near-perfect images. Image denoising is essential in many real-life problems. For example, denoising in MRI, CT scan, and X-ray images leads to clearer images and more accurate diagnoses, In photography, noise removal enhances the clarity of photos captured in dark conditions or with high ISO settings. In astronomy, efficient noise removal aids in the identification of celestial bodies and phenomena. Satellite imaging benefits from denoising by refining images captured from space, allowing better analysis and interpretation of geographic and environmental data. Figure 1 shows the benefits of efficient noise removal from images.
@@
Figure 1: Example image restoration of real world images by Dong et al. (2013).
@@
An overview of the long history of image denoising can be found in Gonzalez and Woods (2018); Qiu (2005). One major type of image denoising algorithms does not detect edges explicitly, but uses the edge information from the image intensities. Bilateral filtering methods e.g., Chu et al. (1998) use the edge information to assign weights in the local intensity averaging procedure. An anisotropic diffusion method proposed by Perona and Malik (1990) uses the edge information to control the direction and the amount of local smoothing. The major drawback of these approaches is that the edges become blurred while smoothing, especially around complicated edge structures such as corners. This is because all pixels in a neighborhood get non-zero weights in the process of local smoothing. Markov random field (MRF) based denoising methods Geman and Geman (1984); Godtliebsen and Sebastiani (1994) introduce a line process to use the edge information, whereas methods such as Rudin et al. (1992); Keeling (2003); Wang and Zhou (2006), etc. introduce various penalty terms in appropriate cost minimization problems. These approaches blur the fine image details to some extent. Other popular types of image denoising methods include wavelet transformation methods Portilla et al. (2003), non-local means algorithms Buades et al. (2005, 2010); Liu et al. (2008), pointwise shape adaptive methods Foi et al. (2007), channel or orientation space methods Felsberg et al. (2006); Franken and Duits (2009) to name a few. Another major type methods explicitly detects the edges first, and then use the information to denoise Qiu (1998); Mukherjee and Qiu (2011). See Qiu (2007) for a detailed description. The explicit edge detection based methods do not work well on various types of images such as low resolution images, textured images, etc. To resolve this issue, Mukherjee and Qiu (2015) proposes local pixel clustering based image denoising where two different regions within a local neighborhood are detected by pixel clustering based on image intensity values. An adaptive smoothing method using various shapes and sizes of the neighborhood for local smoothing depending on local edge information is another approach Takeda et al. (2007). Even though the neighborhoods are elongated along the edges, they often contain pixels on both sides of edges, resulting in image blurring. In spite of having all these methods, no one method outperforms others on a large class on images. It is to be noted that each of these methods has its specific weaknesses alongside its strengths.
@@
Several types of modern image analyses such as image comparisons Roy and Mukherjee (2024a), image monitoring Roy and Mukherjee (2024b) require feature preserving image denoising as a preliminary step, or several such state-of-the-art techniques use the central ideas of various image smoothing methods as building blocks that suit their own specific requirements. Performance of image comparisons and monitoring by such methods depends heavily on the performance of image smoothing. Therefore, it is an urgent requirement that integration of various types of state-of-the-art image denoising methods are done to combine the strengths and remove the weaknesses of the concerned methods as much as possible. Another important aspect to note in this context is that deep learning based denoising methods (e.g., Zamir et al. (2022)) require a large set of images for training or learning. Therefore, such denoising methods cannot be easily integrated with image monitoring procedures which itself is a computationally heavy problem. Therefore, requirement of developing integrated denoising methods such as the one we propose is still exists very much.
@@
The objective of this paper is to integrate the strengths of local clustering based methods and those adaptive methods which use neighborhoods of various shapes and sizes. In this way, complicated edge structures and fine image details should be preserved while removing noise from an image. The steering kernel method by Takeda et al. (2007) choose small neighborhood near complicated edge structures but gives positive weights to all pixels in the neighborhood. In the proposed method, we use a bigger neighborhood in such regions as well and preserve the complicated edge structures by employing local clustering in that neighborhood and averaging the intensities only from a meaningful partition of that neighborhood Mukherjee and Qiu (2015). The proposed method utilizes explicit edge detection to determine neighborhood shapes and sizes specifically for pixels located far from the edges. For pixels close to edges, we employ local clustering methods to achieve enhanced performance, particularly around complicated edge structures. For pixels far away from the edges, we use a large enough elliptical neighborhood to smooth the region as much as possible. Numerical studies show that the practical performance of the proposed method is superior to both these approaches as well as most state-of-the-art denoising methods.
@@ 2 Proposed Methodology
Assume that the observed image intensities following the following jump regression model
$z_{ij}=f(x_{i},y_{i})+\epsilon_{ij}\hskip 7.22743ptfor\hskip 7.22743pti,j=1,2,\dots,n,$
where $\{(x_{i},y_{j})=(i/n,j/n),i,j=1,2,\dots,n\}$ are equally spaced design points or pixels. $\{\epsilon_{i,j}\}$ are independent and identically distributed (i.i.d.) random variables with mean $\mu$ and variance $\sigma^{2}$. $f(x,y)$ is continuous over $[0,1]\times[0,1]$ except on some curves which we call the jump location curve (JLC). A point (i,j) is called a singular point if it satisfies at least one of the following conditions:
  • There exists a constant $\zeta_{0}>0$ such that for any $0<\zeta<\zeta_{0}$ the neighborhood of $(i/n,j/n)$ of diameter $\zeta$ is divided into more than two connected regions by the JLCs.
@@ 2.1 Detection of edge pixels
and we fit a least square plane in this neighborhood:
$\widehat{z}_{ij}(x,y)=\widehat{\beta}_{0}^{(i,j)}+\widehat{\beta}_{1}^{(i,j)}(x-x_{i})+\widehat{\beta}_{2}^{(i,j)}(y-y_{j}),\hskip 7.22743pt(x,y)\in N(x_{i},y_{j})$
@@
$\delta_{ij}=\min\Bigg{(}||\vec{v}_{ij}-\vec{v}_{B_{1}}||,||\vec{v}_{ij}-\vec{v}_{B_{2}}||\Bigg{)},\hskip 7.22743pt\text{where }\vec{v}_{ij}=\left(\widehat{\beta}_{1}^{(i,j)},\widehat{\beta}_{2}^{(i,j)}\right).$
This approach is similar to the one proposed by Qiu and Yandell (1997). $\delta_{ij}$ should be large if and only if $(x_{i},y_{j})$ is close to a JLC. Qiu and Yandell (1997) show that under certain regularity assumptions, $\delta_{ij}>\widehat{\sigma}\sqrt{\frac{\chi^{2}_{2,\alpha_{n}}}{kS_{x}^{2}}}$ can be used as an edge pixel detection criterion. Here, $k$ is the window width and $\chi^{2}_{2,\alpha_{n}}$ is the $(1-\alpha_{n})$ quantile of $\chi^{2}_{2}$ distribution, and $S_{x}^{2}$ is the sample variance of the $X$-coordinates in a local neighborhood.
@@ 2.2 Local smoothing
In this step of the proposed method, we find an elliptical neighborhood around the each pixel of the image such that no detected edge pixels are inside it. To accomplish this, we execute the following steps:
@@
    Next we consider the point $M$, which is the reflection of the point $P_{1}$ with respect to the given point. Then, we find the nearest edge pixel in the strip perpendicular to the line joining $M$ and $P_{1}$, say $P_{2}$.
@@
    Now, consider the ellipse $E$, which is centered at $P$ and has major axis of length $|PP_{2}|-\gamma$ and minor axis of length $|PP_{1}|-\gamma$. Observe, we also have $d(P_{1},E)>\gamma$ and $d(P_{2},E)>\gamma$. Here $\gamma$ is a tuning parameter. See Figure 2 for demonstration.
@@
where $f$ is an unknown function and $\tilde{p}$ is the coordinates of a pixel on the image. We assume that the true intensity function $f$ is smooth up to $N$-th order. If $\tilde{p}_{i}$ is close to $\tilde{p}$, we can use the Taylor series expansion to get:
$z(\tilde{p}_{i})$ $\approx z(\tilde{p})+(\nabla z(\tilde{p}))^{T}(\tilde{p}_{i}-\tilde{p})+\frac{1}{2}(\tilde{p}_{i}-\tilde{p})^{T}\mathcal{H}_{z}(p)(\tilde{p}_{i}-\tilde{p})+\dots$
@@
where $\theta_{0}=z(\tilde{p})$, $\theta_{1}=\nabla z(\tilde{p})$, $\theta_{2}=\frac{1}{2}\Bigg{[}\frac{\partial^{2}z(\tilde{p})}{\partial\tilde{p}_{1}^{2}},2\frac{\partial^{2}z(\tilde{p})}{\partial\tilde{p}_{1}\partial\tilde{p}_{2}},\frac{\partial^{2}z(\tilde{p})}{\partial\tilde{p}_{2}^{2}}\Bigg{]}$, $vech\begin{pmatrix}a&b\\ b&d\end{pmatrix}=(a,b,d)$ and $\mathcal{H}_{z}(p)$ is the Hessian matrix evaluated at $\tilde{p}$.
Therefore, we formulate this as a regression problem and estimate $\theta_{0}$, $\theta_{1}$, $\dots$ by minimizing :
$\sum_{i=1}^{N_{p}}\bigg{[}z(\tilde{p}_{i})-\theta_{0}-\theta_{1}^{T}(\tilde{p}_{i}-\tilde{p})-\theta_{2}^{T}vech{(\tilde{p}_{i}-\tilde{p})(\tilde{p}_{i}-\tilde{p})^{T}}+...\bigg{]}^{2}K\left(\frac{\tilde{p}_{i}-\tilde{p}}{h}\right)$
@@
For pixels which are closer to the detected edge pixels, i.e., within $\gamma$ distance of an edge pixel, we are using the clustering method proposed by Mukherjee and Qiu (2015). Consider a neighborhood $O(x,y,h_{n})$ around one such point $(x,y)$. We can choose a threshold $s$, so that the pixels with intensity value $z_{ij}\leq s$ are i one cluster. Then we are choosing s such that it minimizes $T(x,y,h_{n},s)$, the ratio of between-group and within-group variability of the intensity values.
@@
where $O_{1}(x,y,h_{n},s)$ and $O_{2}(x,y,h_{n},s)$ are the two clusters created using threshold $s$. And $\bar{z}_{1}$ and $\bar{z}_{2}$ are the average of intensity values in the respective clusters. We can choose the optimal value of $s$ using:
@@
After dividing the neighborhood $O(x,y,h_{n})$ into to clusters, we estimate the true image intensity at $(x,y)$ using the weighted average of intensity values of the pixels in the cluster containing $(x,y)$. In this paper, we use a similarity measure between pixels as weights, which is given by :
$\tilde{W}_{ij}=exp\Bigg{(}-\frac{||\tilde{O}(x_{i},y_{j})-\tilde{O}(x,y)||^{2}_{2}}{2\hat{\sigma}^{2}|\tilde{O}(x_{i},y_{j})|B_{n}}\Bigg{)},$
where $\tilde{O}(x,y)$ is a neighborhood around $(x,y)$ of radius $\tilde{h}_{n}$, $\hat{\sigma}$ is the estimated standard deviation of errors. and $B_{n}$ is a tuning parameter. We are using $L_{2}$ distance between the observed intensity values as the distance here.
To summarize, our algorithm can be represented by the flow chart below. Please read the chart along with Figure 2.
Pick a pixel $P$ from the image Find the nearest edge pixel $P_{1}$ and calculate the distance between $P$ and $P_{1}$ $\text{Distance}(P,P_{1})<\gamma$ Find nearest pixel $P_{2}$ in the strip perpendicular $P_{1}M$ where $M$ is reflection of $P_{1}$ with respect to $P$. no Smooth using clustering method yes Construct an ellipse with $P$ as center, and $|PP_{1}|-\gamma$,$|PP_{2}|-\gamma$ as length of the axes Smooth in the neighborhood E around P using kernel regression.

### v3

@@ Abstract
The importance of developing efficient image denoising methods is immense especially for modern applications such as image comparisons, image monitoring, medical image diagnostics, and so forth. Available methods in the vast literature on image denoising can address certain issues in image denoising, but no one single method can solve all such issues. For example, jump regression based methods can preserve linear edges well, but cannot preserve many other fine details of an image. On the other hand, local clustering based methods can preserve fine edge structures, but cannot perform well in presence of heavy noise. The proposed method uses various shapes and sizes of local neighborhood based on local information, and integrates this adaptive approach with the local clustering based smoothing. Theoretical justifications and numerical studies show that the proposed method indeed performs better than these two individual methods, and outperforms many other state-of-the-art techniques as well. This superior performance emphasizes the strong potential of the proposed method for broad applicability in modern image analysis.
Keywords: Adaptive smoothing, Edge preservation, Local clustering, Multi-resolution smoothing, Variable neighborhood.
@@ 1 Introduction
Easy access to modern image acquisition techniques is making images a popular data format in various disciplines of science. Applications in manufacturing industries include stress and strain analysis of products, anomaly detection in the rolling process, steel and tile surface monitoring, etc. In medical science, various imaging modalities such as X-ray, CT-scan, MRI, and fMRI are being widely used for medical diagnosis. For surveillance of the Earth’s surface, satellite images are commonly used. These images serve as a fundamental tool for research in agriculture, forest science, ecology, ecosystems, and many more. Note that image comparisons and sequential image monitoring are important in all of these applications. These images often contain noise, anomalies, etc. that make eventual image analysis unreliable. Therefore, image analysis as a modern-day research area is increasingly gaining a lot of importance. See Qiu (2018) for related discussions.
Along with increasing popularity of digital imaging systems, there is a growing demand for near-perfect images. Image denoising is essential in many real-life problems. For example, denoising techniques applied to MRI, CT scan, and X-ray images significantly improve image quality, thereby facilitating more accurate clinical diagnoses. In photography, noise removal enhances the clarity of photographs captured in dark conditions or with high ISO settings. In astronomy, efficient noise removal aids in the identification of celestial bodies and phenomena. Satellite imaging benefits from denoising by refining images captured from space, allowing better analysis and interpretation of geographic and environmental data. Figure 1 shows the benefits of efficient noise removal from images.
@@
Figure 1: Image restoration of real world images provided by Dong et al. (2013).
@@
An overview of the long history of image denoising can be found in Gonzalez and Woods (2018); Qiu (2005). One major type of image denoising algorithms does not detect edges explicitly, but uses the edge information from the observed image intensities. Bilateral filtering methods e.g., Chu et al. (1998) use the edge information to assign weights in the local intensity averaging procedure. The anisotropic diffusion method proposed by Perona and Malik (1990) uses the edge information to control the direction and the amount of local smoothing. The major drawback of these approaches is that the edges become blurred while smoothing, especially around complicated edge structures such as corners. This is because all pixels in a neighborhood get non-zero weights in the process of local smoothing. Markov random field (MRF) based denoising methods Geman and Geman (1984); Godtliebsen and Sebastiani (1994) introduce a line process to use the edge information, whereas methods such as Rudin et al. (1992); Keeling (2003); Wang and Zhou (2006), etc. introduce various penalty terms in appropriate cost minimization problems. These approaches blur the fine image details to some extent. Other popular types of image denoising methods include wavelet transformation methods Portilla et al. (2003), non-local means algorithms Buades et al. (2005, 2010); Liu et al. (2008), pointwise shape adaptive methods Foi et al. (2007), channel or orientation space methods Felsberg et al. (2006); Franken and Duits (2009) to name a few. One major approach to solve the problem of image denoising is through jump regression analysis Qiu (2005). The majority of the initial such attempts Qiu (1997, 1998) detect the edges first, and then use these edge information to denoise the image. Later, Mukherjee and Qiu (2011) extend similar ideas to denoise 3-D images. Jump regression based image denoising techniques that do not explicitly detect edges include methods by Qiu (2004); Gijbels et al. (2006). See Qiu (2007) for a detailed discussion. The explicit edge detection based methods do not work well on various types of images such as low resolution images, textured images, etc. To resolve this issue, Mukherjee and Qiu (2015) propose local pixel clustering based image denoising where two different regions within a local neighborhood are detected by pixel clustering based on image intensity values. An adaptive smoothing method using various shapes and sizes of the neighborhood for local smoothing depending on local edge information is another approach (e.g., Takeda et al. (2007)). Even though the neighborhoods are elongated along the edges, they often contain pixels on both sides of edges, resulting in image blurring. In spite of having all these methods, no one method outperforms others on a large class on images. It is to be noted that each of these methods has its specific weaknesses alongside its strengths.
@@
Several types of modern image analyses such as image comparisons Feng and Qiu (2018); Roy and Mukherjee (2024); Guha and Qiu (2025), image monitoring Roy and Mukherjee (2024) require feature preserving image denoising as a preliminary step. Sometimes, many such methods use the central ideas of various image smoothing methods as building blocks that suit their own specific requirements. Performance of image comparison and monitoring by such methods depends heavily on the effectiveness of image smoothing. Therefore, it is an urgent requirement that integration of various types of state-of-the-art image denoising methods are done to combine the strengths and remove the weaknesses of the concerned methods as much as possible. Another important aspect to note in this context is that deep learning based denoising methods (e.g., Zamir et al. (2022)) require a large set of images for training or learning. Therefore, such type of denoising methods cannot be easily integrated with image monitoring procedures which itself is a computationally heavy problem. Therefore, requirement of developing integrated denoising methods such as the one we propose still exists very much.
@@
The objective of this paper is to integrate the strengths of local clustering based methods and those adaptive methods which use neighborhoods of various shapes and sizes. In this way, complicated edge structures and fine image details should be preserved while removing noise from an image. The steering kernel method by Takeda et al. (2007) choose small neighborhood near complicated edge structures but gives positive weights to all pixels in the neighborhood. In the proposed method, we use a bigger neighborhood in such regions as well, and preserve the complicated edge structures by employing local clustering in that neighborhood and averaging the intensities only from a meaningful partition of that neighborhood Mukherjee and Qiu (2015). The proposed method utilizes explicit edge detection to determine neighborhood shapes and sizes. For pixels close to the edges, we employ local clustering methods to achieve enhanced performance, particularly around complicated edge structures. For pixels far away from the edges, we use a large enough elliptical neighborhood to smooth the region as much as possible. Numerical studies show that the practical performance of the proposed method is superior to both these approaches as well as most state-of-the-art denoising methods.
@@ 2 Proposed Methodology
Assume that the observed image intensities follow the jump regression model
$z_{ij}=f(x_{i},y_{i})+\epsilon_{ij}\hskip 7.22743pt\mbox{for}\hskip 7.22743pti,j=1,2,\dots,n,$
where $\{(x_{i},y_{j})=(i/n,j/n),\ i,j=1,2,\dots,n\}$ are equally spaced design points or pixels. $\{\epsilon_{i,j}\}$ are independent and identically distributed (i.i.d.) random variables with mean 0, and variance $\sigma^{2}$. We assume that $f(x,y)$ is continuous over $0,1]\times[0,1]$ except on some curves which we call the jump location curve (JLC). A point (i,j) is called a singular point if it satisfies at least one of the following conditions:
  • There exists a constant $\zeta_{0}>0$ such that for any $0<\zeta<\zeta_{0}$, the neighborhood of $(i/n,j/n)$ of diameter $\zeta$ is divided into more than two connected regions by the JLCs.
@@ 2.1 Detection of edge pixels
and we fit a least square plane in this neighborhood around $(x_{i},y_{j})$:
$\widehat{z}_{ij}(x,y)=\widehat{\beta}_{0}^{(i,j)}+\widehat{\beta}_{1}^{(i,j)}(x-x_{i})+\widehat{\beta}_{2}^{(i,j)}(y-y_{j}),\hskip 7.22743pt(x,y)\in N(x_{i},y_{j}).$
@@
[…]
This approach is similar to the one proposed by Qiu and Yandell (1997). $\delta_{ij}$ should be large if and only if $(x_{i},y_{j})$ is close to a JLC. Qiu and Yandell (1997) show that under certain regularity assumptions, $\delta_{ij}>\widehat{\sigma}\sqrt{\frac{\chi^{2}_{2,\alpha_{n}}}{kS_{x}^{2}}}$, which can be used as an edge pixel detection criterion. Here, $k$ is the window width, $\chi^{2}_{2,\alpha_{n}}$ is the $(1-\alpha_{n})$ quantile of $\chi^{2}_{2}$ distribution, and $S_{x}^{2}$ is the sample variance of the $X$-coordinates in the local neighborhood.
@@ 2.2 Local smoothing
In this step of the proposed method, we find an elliptical neighborhood around each pixel of the image such that no detected edge pixels are inside it. To accomplish this, we execute the following steps:
@@
    Next, we consider the point $M$, which is the reflection of the point $P_{1}$ with respect to the given point. Then, we find the nearest edge pixel from $P$ in the strip perpendicular to the line joining $M$ and $P_{1}$. Call that edge pixel $P_{2}$. See Figure 2 for visual explanation.
@@
    Now, consider the ellipse $E$, centered at $P$, with half-lengths of the major and minor axes given by $|PP_{2}|-\gamma$ and $|PP_{1}|-\gamma$, respectively. The tuning parameter $\gamma$ ensures that the selected ellipse neither intersects nor touches any JLC.
The point on the ellipse $E$ along the minor axis that is closest to $P_{1}$ corresponds to the intersection of the line segment $PP_{1}$ and $E$. This intersection point, denoted by $P^{\ast}$, lies exactly $\gamma$ distance away from $P_{1}$, i.e., $|P^{\ast}P_{1}|=\gamma$. If a point $\overline{P}$ on a JLC lies within a distance less than $\gamma$ from the ellipse, then, by the triangle inequality, it follows that $|P\overline{P}|<|PP_{2}|$, which contradicts the assumption that $P_{2}$ is the nearest point on the JLC to $P$ within the strip. Therefore, the ellipse maintains a minimum distance of $\gamma$ from all points on any JLC. See Figure 2 for demonstration.
@@
where $f$ is an unknown function and $\tilde{p}$ is the vector of coordinates of a pixel on the image. We assume that the true intensity function $f$ is smooth up to $L$-th order. If $\tilde{p}_{i}$ is close to $\tilde{p}$, we can use the Taylor series expansion to get:
[…]
@@
[…]
$\sum_{i=1}^{N_{p}}\bigg{[}z(\tilde{p}_{i})-\theta_{0}-\theta_{1}^{T}(\tilde{p}_{i}-\tilde{p})-\theta_{2}^{T}vech{(\tilde{p}_{i}-\tilde{p})(\tilde{p}_{i}-\tilde{p})^{T}}+...\bigg{]}^{2}K\left(\frac{\tilde{p}_{i}-\tilde{p}}{h}\right),$
@@
For pixels which are closer to the detected edge pixels, i.e., within $\gamma$ distance of an edge pixel, we are using the clustering method proposed by Mukherjee and Qiu (2015). Consider a neighborhood $O(x,y,h_{n})$ around one such point $(x,y)$. We can choose a threshold $s$, so that the pixels with intensity value $z_{ij}\leq s$ are in one cluster. Then, we are choosing $s$ such that it maximizes $T(x,y,h_{n},s)$, the ratio of between-group and within-group variability of the intensity values.
@@
where $O_{1}(x,y,h_{n},s)$ and $O_{2}(x,y,h_{n},s)$ are the two clusters created using threshold $s$, $\bar{z}_{1}$ and $\bar{z}_{2}$ are the average of intensity values in the respective clusters. We can choose the optimal value of $s$ using:
@@
After dividing the neighborhood $O(x,y,h_{n})$ into two clusters, we estimate the true image intensity at $(x,y)$ using the weighted average of intensity values of the pixels in the cluster containing $(x,y)$. In this paper, we use a similarity measure between pixels as weights, which is given by:
[…]
where $\tilde{O}(x_{i},y_{j})-\tilde{O}(x,y)||^{2}_{2}$ is the sum of squares of the pixelwise differences of the intensities between the neighborhoods $\tilde{O}(x_{i},y_{j})$ and $\tilde{O}(x,y)$, $|\tilde{O}(x_{i},y_{j})|$ is the number of pixels in the neighborhood $\tilde{O}(x_{i},y_{j})$, $\hat{\sigma}$ is the estimated standard deviation of errors, and $B_{n}$ is a tuning parameter for smoothing.
The following flow chart summarizes the proposed integrated image denoising method. Please read the chart along with Figure 2.
Pick a pixel $P$ from the image. Find the nearest edge pixel $P_{1}$, and calculate $|PP_{1}|$. $|PP_{1}|<\gamma$ Find nearest pixel $P_{2}$ in the strip perpendicular to $P_{1}M$, where $M$ is reflection of $P_{1}$ with respect to $P$. no Smooth using clustering method. yes Construct an ellipse with $P$ as center, and $|PP_{1}|-\gamma$, $|PP_{2}|-\gamma$ as half-lengths of the axes. Smooth in the neighborhood E around P using kernel regression.